OpenStudy (anonymous):

The polygons are similar, but no necessarily drawn to scale. Find the value of x. Show your work.

OpenStudy (bibby):

If you notice, all the sides are being multiplied by a factor of 4. that means that 4*[that side] = x-3

OpenStudy (bibby):

4*2.5 = 10 10 = x-3

OpenStudy (bibby):

if the whole side is 10, and the side is x-3, what is x?

OpenStudy (bibby):

Sorry, I'm being vague. I'll try and re explain it. \[\huge 4*2.5 = (x-3)\] that x-3 is the whole side. You can just solve for x here

OpenStudy (bibby):

totes wrong Start by multiplying 2,5 and and 4, what does that give you?

OpenStudy (bibby):

10 = x - 3, solve for x

OpenStudy (bibby):

You've solved for x, hooray

OpenStudy (bibby):

btw in case you don't see how I got the 4, you set up a proportion like so:

OpenStudy (bibby):

You either look at the ratio of 16:4 and divide that, till you get 4:1 or you can use this: \[\huge \frac{ 16 }{ 4 } = \frac{ x-3 }{ 2.5 }\]

OpenStudy (bibby):

Yeah post another question and tag me there